### v1

# M3GIA: A Cognition Inspired Multilingual and Multimodal General Intelligence Ability Benchmark

Abstract
As recent multi-modality large language models (MLLMs) have shown formidable proficiency on various complex tasks, there has been increasing attention on debating whether these models could eventually mirror human intelligence. However, existing benchmarks mainly focus on evaluating solely on task performance, such as the accuracy of identifying the attribute of an object. Combining well-developed cognitive science to understand the intelligence of MLLMs beyond superficial achievements remains largely unexplored. To this end, we introduce the first cognitive-driven multi-lingual and multi-modal benchmark to evaluate the general intelligence ability of MLLMs, dubbed M3GIA. Specifically, we identify five key cognitive factors based on the well-recognized Cattell-Horn-Carrol (CHC) model of intelligence and propose a novel evaluation metric. In addition, since most MLLMs are trained to perform in different languages, a natural question arises: is language a key factor influencing the cognitive ability of MLLMs? As such, we go beyond English to encompass other languages based on their popularity, including Chinese, French, Spanish, Portuguese and Korean, to construct our M3GIA. We make sure all the data relevant to the cultural backgrounds are collected from their native context to avoid English-centric bias. We collected a significant corpus of data from human participants, revealing that the most advanced MLLM reaches the lower boundary of human intelligence in English. Yet, there remains a pronounced disparity in the other five languages assessed. We also reveals an interesting winner takes all phenomenon that are aligned with the discovery in cognitive studies. Our benchmark will be open-sourced, with the aspiration of facilitating the enhancement of cognitive capabilities in MLLMs.
1 Introduction
In 1956, researchers across different domains, including mathematics, cognitive psychology and computer science, pointed out an interesting direction, dubbed artificial intelligence (AI). The formal definition is “The study is to proceed on the basis of the conjecture that every aspect of learning or any other feature of intelligence can in principle be so precisely described that a machine can be made to simulate it.” [28]. Through extensive efforts in pursuing artificial intelligence, the field has converged to a paradigm of data-driven machine learning models, which are still deeply intertwined with cognitive science as they often mirror basic cognitive mechanisms, e.g. convolutional neural networks [18] and the attention mechanism [43]. Recent advances, such as GPT-4o [30], demonstrate that these MLLMs can outperform human on various complex tasks [1, 45] and shed light to emergent ability with the increasing scale of data and model size [46]. In light of these developments, our aim is to evaluate these state-of-the-art models through the lens of cognitive science, as it directly aligns with the primary motivation of AI research.
To explore the mental intelligence emerging from these large models, efforts have been directed toward analyzing these models from a psychological perspective. Some pioneering works report that LLMs have demonstrated human-like cognition [4, 17]. For instance, Theory of mind (ToM) has been applied to assess large models, revealing that GPT-4 exhibits ToM capabilities similar to human inference patterns [5, 17, 15]. Meanwhile, Multimodal Large Language Models (MLLMs), which use powerful LLMs as brain to process and integrate multimodal information, have exhibited impressive emergent abilities, such as generating website code from images [56], understanding the meaning of a meme [48], and math reasoning [12]. Thanks to their ability to process information from a broader spectrum of sources, they exhibit a more holistic cognitive process, resembling human cognition more closely than models confined to purely linguistic input.
Figure 1: Overview of multi-lingual multi-modal general intelligence ability benchmark. (Left) In contrast to traditional benchmarks that focus on evaluating specific task performances, we draw inspiration from cognitive science to categorize five cognitive factors, try to provide a feasible evaluation of general intelligence ability (GIA). (Right) Specifically, we adopt the factors from the CHC theory to disentangle fundamental cognitive abilities with existing evaluation tasks. In addition, to further understand how language impacts such ability, we collect or design questions in six languages with large population.
Existing multi-modality benchmarks, such as MMBench [23], MME [14], and MM-Vet [52], have made the attempt to compartmentalize model capabilities across multiple tasks. For instance, MMBench covers 20 different abilities, encompassing function reasoning, physical property reasoning, object localization and social reasoning. However, they often fail to provide a persuasive explanation for their selection of dimensions, as they tend to be mired in subjectivity and lack a solid theoretical underpinning. Moreover, as depicted in Figure 1 (left), their ability dimensions are still rather task-oriented, neglecting a systematic evaluation of the models’ underlying cognitive abilities that govern task performance through the lens of cognitive science. This oversight raises concerns that benchmarks might devolve into mere training targets rather than instruments for true insight, failing to provide a holistic measure of the models’ capabilities [35]. In short, the ability to solve specific tasks is insufficient to reflect the true level of intelligence, as supported by a psychological study[31], and formally evaluating the cognitive factors of MLLMs remains largely unexplored.
In this paper, we close the gap by introducing the first benchmark that comprehensively evaluate the cognitive abilities of MLLMs under the theoretical umbrella of the well-recognized Cattell-Horn-Carroll (CHC) Model of Intelligence [36], dubbed M3GIA. As in Figure 1(right), based on the CHC Model, we categorizes the cognitive capacities of current MLLMs into five dimensions: Fluid reasoning (Gf), Comprehension-Knowledge (Gc), Visual processing (Gv), Reading and Writing (Grw), Quantitative knowledge (Gq), and collect corresponding questions as a measurement. In addition, as using multi-lingual data to scale up the capability of MLLMs becomes a de-facto standard, we are curious whether languages make any impact on their cognitive abilities. As such, we extend our benchmark to include five more languages, including Chinese, Spanish, French, Portuguese and Korean roughly based on their population, to disentangle the language factor with cognitive ability.
To evenly assess the five cognitive dimensions, we refer to human intelligence tests, such as Raven’s Progressive Matrices Test [32] and the Woodcock-Johnson IV Tests of Cognitive Abilities (WJ IV) [38], and establish broad question types that correspond to the these cognitive dimensions, which are further subdivided into 18 narrow question types (see later Sec. 3). All in all, our M3GIA contains 1,800 questions, where over half are carefully designed from scratch following the standard. The test for each language maintain consistency in terms of the number of questions, structure, and distribution of question types. In addition, to highlight the multilingual nature of our benchmark, we collect data relevant to cultural backgrounds from native language sources rather than simply translating them from English, thereby avoiding the English-centric bias.
We evaluate 24 MLLMs, including the state-of-the-art close and open-sourced ones. In general, The latest advancements in MLLMs have achieved performance levels that fall within the lower boundary of human intelligence in English. Yet, there remains a pronounced disparity in the other five languages assessed. We also notice that MLLMs’ proficiency in one cognitive domain often translates into superior performance across other domains as well. This phenomenon interestingly aligns with the pattern observed in human intelligence which empirically suggests the existence of General Intelligence Ability in MLLMs.
2 Related works
Evaluation Benchmark for MLLMs.
As multimodal large language models (MLLMs) exhibit remarkable generalization capabilities across a broad spectrum of downstream tasks, relying exclusively on their performance within single vision-language tasks — such as visual recognition [16], image description [8, 2, 51], scene text understanding [40, 39], and external knowledge [27] — is insufficient to fully uncover the comprehensive performance of MLLMs. People then turn to a new paradigm to construct all-round benchmarks to assess a broader spectrum of challenging multimodal tasks [49, 47, 20, 23, 14, 52]. Another trend in MLLM assessment is the use of human exam questions [26, 25, 55, 53, 54]. For instance, AGIEval [55] sources questions from standardized exams such as college entrance exams and lawyer qualification tests. While these benchmarks makes progresses in evaluating the human-centric ability of MLLMs, it may not be suitable to evaluate the intelligence of MLLMs because research in psychological field points out that the superficial performance on tasks alone cannot be a solid indicator for human’s intelligence.[31]
General Intelligence Ability and the CHC Theory.
Arising from the empirical fact that an individual’s proficiency in one area frequently correlates with high performance in other areas, Charles Spearman first introduced General Intelligence Ability (GIA) in 1904 [41]. This construct refers to the idea that a single underlying factor, known as the g-factor, can account for the positive correlations among cognitive abilities and reflect the general intelligence that fundamentally underlies an individual’s intelligence. To concretely understand GIA, numerous attempts has been made to model the structure of human cognition. John Carroll’s Three-Stratum Model [7] elaborated on this with a hierarchical structure of intelligence, including a general “g” factor and specific cognitive abilities. Howard Gardner’s Multiple Intelligence Theory [13] proposed diverse forms of intelligence, while Sternberg’s Triarchic Theory [42] focused on practical, creative, and analytical aspects. These theories collectively contributed to the development of the Cattell-Horn-Carroll (CHC) model of intelligence, which is the most comprehensive and empirically validated structural model of human cognition [29] to date, integrating various aspects of cognition into a unified framework. Recent study [11], has primarily focused on evaluating the performance of large language models on sophisticated psychological tasks, neglecting the assessment of models’ intelligence from the foundational standpoint of cognitive models. Our M3GIA constitutes the first attempt to bring the latest cognitive science modeling of intelligence into MLLMs evaluation to address this gap.
3 M3GIA
Concretely, we introduce the first cognition inspired multi-linguistic and multi-modal benchmark to evaluate the general intelligence accuracy of large models. In short, our M3GIA distinguishes itself from existing benchmarks as follow:
  • Cognition Inspired: In contrast to existing benchmarks that focuses on task-level evaluation, we study the intelligence of large models from a cognition perspective. The benchmark dissects the cognitive abilities of contemporary MLLMs into five foundational factors, as per the Cattell-Horn-Carroll theory. This cognitive theory underpins the structure of our evaluation, informing the specific types of questions devised to test each cognitive skill.
  • Multilingual Coverage: To comprehensively measure the cognitive abilities of multimodal large models across multiple languages, M3GIA is constructed to span six languages: English, French, Chinese, Spanish, Portuguese, and Korean. In order to mitigate English-centric bias, all data relevant to cultural backgrounds have been sourced from native language resources, except for questions that transcend cultural considerations—such as the Raven test and number series problems.
The subsequent content of this section is organized as follows: In sec. 3.1, we introduce the five-factor cognitive model of M3GIA and discuss the design philosophy behind it. In sec. 3.2, we describe how we designed and collected the questions for M3GIA and provide some statistical data on M3GIA.
Figure 2: Structure of our CHC inspired model of cognitive abilities. (Left) We identified five key cognitive factors for current MLLMs: Comprehension-Knowledge (Gc), Quantitative knowledge (Gq), Reading and Writing (Grw), Fluid reasoning (Gf), and Visual-spatial processing (Gv). In the hierarchical structure, Gf is further subdivided into three narrow factors: I (Induction), RG (Deductive Reasoning), and RQ (Quantitative Reasoning). (Right) A conceptual map of the five cognitive factors and their overlaps with each other.
3.1 The Five-factor Cognitive Model of M3GIA
To formally study the large models intelligence level, we start from the state-of-the-art cognitive model, Cattell-Horn-Carroll (CHC) [36], which is by far the most empirically validated structure model of human cognition [29]. The CHC theory articulates a hierarchical framework of human cognitive abilities divided into three strata: general intelligence “g” (stratum III), broad cognitive abilities (stratum II), and narrow cognitive abilities (stratum I). The theory has now expanded to include 16 broad abilities and over 80 narrow abilities. These broad but domain-specific abilities are nevertheless positively associated with one another. This positive manifold is accounted for in the CHC model by a general factor of intelligence (“g”) at stratum III. While there is ongoing discourse regarding the exact delineation of the narrow abilities, 9 out of the 16 broad cognitive abilities have achieved substantial consensus and are well-supported by empirical evidence and practical application.[6] These include Fluid Reasoning (Gf), Comprehension-Knowledge (Gc), Visual Processing (Gv), Auditory Processing (Ga), Short-term Memory (Gsm), Long-term Retrieval (Glr), Processing Speed (Gs), Quantitative Knowledge (Gq), and Reading and Writing Abilities (Grw).
Figure 3: Questions overview of M3GIA. To assess the five CHC cognitive factors—Gf, Gc, Gq, Grw, Gv correspondingly—we devised five broad question clusters: common sense (orange), visual-spatial (blue), comprehension (yellow), mathematics (red), and reasoning (green). To prevent the assessment of any particular ability from being constrained to a fixed and singular perspective, we have stratified each of the five clusters into 2-4 specialized narrow question types, each narrow question type reflects a different perspective on the broad CHC ability. This subdivision results in a total of 18 subtasks. All the QAs are in the format of multiple choice problems whose answers are marked [A][B][C][D].
As shown in Fig. 2, the structure of our M3GIA is underpinned by the five-factor hierarchical cognitive model, which is derived from the CHC model of cognitive abilities. Although Large Language Models exhibit cognitive processes similar to humans, they also differ in internal mechanisms, particularly with regard to processing speed (Gs) and memory (Gwm, Glr), which is greatly related to external technologies beyond the model itself, such as external databases and retrieval-augmented generation (RAG) [19]. Additionally, given that the majority of current MLLMs, with the exception of a select few closed-source models, are not yet expanded to embrace the auditory modality, we have not included the Ga (Auditory Processing) factor in this version of M3GIA, reserving it as one of the directions for future expansion. Consequently, based on the consultations with psychology experts, we have chosen to assess the cognitive abilities of current MLLMs in this iteration of M3GIA by focusing on five key CHC factors: Gc, Grw, Gq, Gf, and Gv, from among the nine most frequently identified CHC factors.
Interestingly, the five factors we select align closely with those of the renowned Stanford-Binet Test, Fifth Edition (SB5) [34], which was also constructed upon five cognitive factors derived from the CHC theory. Specifically, the five cognitive factors identified in the SB5 are: Fluid Reasoning (FR), Knowledge (KN), Quantitative Reasoning (QR), Visual-Spatial Processing (VS), and Working Memory (WM). Except for Working Memory (WM), which we have substituted with Grw, these factors align directly with our selected factors, corresponding to Gf, Gc, Gq, and Gv, respectively. This alignment is noteworthy, as the selection of these factors for the SB5 was based on extensive research on school achievement and expert ratings of the importance of these factors in the assessment of reasoning, especially in giftedness assessment [33].
3.2 Question Design and Collection
Our M3GIA contains a total of 1,800 multiple choice problems, of which 1,200 are Visual Question Answering (VQA) questions. The ground truths for 630 questions are human-annotated, while the remainder of the answers for 1,170 questions were gathered from the Internet.
As shown in Fig. 1, we have devised five broad question clusters: reasoning, visual-spatial, common sense, mathematics and comprehension, separately corresponding to the assessment of the five CHC cognitive factors – Gf, Gv, Gc, Gq, and Grw. See supplementary for more details. To prevent the assessment of any particular ability from being constrained to a fixed and singular perspective, we have stratified each of the five clusters into 2-4 narrow question types that reflect different perspectives on a broad CHC construct. This subdivision results in a total of 18 distinct question types, each designed to tap into different facets of the ability being measured. Consequently, any generalizations that are made from a cluster are based on two or more samples of ability, which reduces the possibility of making over-generalizations from a narrow sampling of ability.
Moreover, as illustrated in the right part of Fig. 2, the five cognitive factors are not isolated but rather overlap with each other. For example, Fluid reasoning (Gf) not only has a process facet (inductive vs. deductive reasoning) but also has a content facet (verbal, spatial, and quantitative), each of which overlaps with other broad abilities. [37]. In order to conduct a comprehensive measurement of this overlapping nature, our narrow question types include not only tests that measure each cognitive factor individually but also cover the parts where these factors overlap. The corresponding relationships between the question types, the cognitive factors and their intersections are also shown in Fig. 2.
What is more, to ensure that our assessment remains anchored in reality, we incorporate real-world problems into the evaluations of cognitive abilities. Specifically, each broad question type includes not only abstract cognitive test questions but also typical real-world problems that require the use of one or more cognitive abilities. This approach enables us to conduct a more accurate and practical assessment of how well these abilities are applied outside controlled, test-like environments. To ensure a balanced and comprehensive evaluation for each ability, we have tried our best to maintain an even distribution among problems associated with different abilities during data collection.
Examples of the narrow question types can be seen in Fig. 3, while more detailed descriptions are included in the Appendix.
3.3 Metrics
We use two type of metrics in our evaluation benchmark. For each narrow question type, we follow the existing benchmarks [23, 14] to use accuracy. However, to holistically compare the cognitive ability, we design a novel metric general intelligence accuracy (GIA) based on findings in cognitive field. To compute the GIA scores of the models and validate the consistency of the cognitive structure between MLLMs and human intelligence, we adopted a standard psychometric approach. This involved utilizing a confirmatory factor analysis (CFA) model, developed from our collected human evaluation data. For more details about the CFA process, see Appendix for more details.
4 Evaluation Results
In this section, we evaluate a total of 24 MLLMs and 480 human participants using our M3GIA. The MLLMs comprise both closed-source models, such as GPT-4o [30], and open-source models [22, 21, 50, 3, 44, 24, 9], including LLaVA [22] and Mini-Gemini [21]. Our evaluation for the MLLMs is conducted under a zero-shot setting to assess the capability of models to generate accurate answers without fine-tuning or few-shot demonstrations on our benchmark. For all models, we conduct prompt engineering on the validation set and use the most effective prompt for the zero-shot setup in the experiments. All experiments are conducted with NVIDIA A800 GPUs [22, 21].
Human Performance Baseline.
To establish a reference for human cognitive levels against MLLMs, we collected 480 valid sets of test data from human subjects using electronic questionnaires. These 480 participants were from native countries of the six selected languages, with 80 individuals per language. The 1,800 questions of M3GIA are then divided into six complete sub-questionnaires by language, with each individual only responsible for completing the sub-questionnaire corresponding to their native language. See supplementary for more details.
4.1 Accuracy Score on Five Cognitive Factors
Table 1: The accuracy results on 24 MLLMs regarding each cognitive ability. The best in bold and the second-best underlined. All the numbers are presented in decimal and the full score is 100.
Types (LLM Size) Models ViT Size Gf Gc Gq Grw Gv Overall Acc
I RG RQ Overall
Human Average Performance - 86.8 60.0 71.2 69.7 79.1 65.4 78.1 81.1 76.9
API GPT-4o - 58.0 59.2 33.9 50.1 72.3 42.8 79.6 46.3 59.8
GPT-4v - 56.7 56.3 40.9 51.9 74.8 46.4 77.5 52.4 59.2
Gemini-1.5-Pro - 54.3 56.4 41.8 54.3 75.8 60.8 77.1 53.8 62.4
Gemini-Pro - 39.0 30.8 22.7 32.4 56.5 31.7 67.1 43.1 46.5
Cluade3-Sonnet - 39.7 32.9 27.3 34.0 58.3 34.2 61.3 43.9 47.0
Cluade3-Haiku - 35.3 35.8 30.3 33.1 55.8 33.3 57.9 36.4 43.1
Mini-Gemini-34b 0.3B 37.7 37.5 30.6 34.8 61.0 34.2 62.9 45.7 48.2
OSS Mini-Gemini-8*7b 0.3B 28.7 30.0 26.7 30.3 58.1 35.0 61.3 41.9 44.8
(Large) LLaVA-v1.6-34b 0.3B 20.7 40.0 28.5 30.8 53.8 36.4 61.7 40.4 42.8
Yi-VL-34b 0.6B 25.0 32.9 35.8 29.5 48.1 29.2 54.6 35.7 38.2
InternVL-chat-v1.2-plus 6B 45.0 42.5 32.4 42.5 64.6 41.4 66.7 47.5 51.9
OSS Mini-Gemini-13b 0.3B 22.3 29.2 23.3 24.3 41.5 26.1 44.2 28.3 32.9
(Medium) LLaVA-v1.5-13b 0.3B 17.7 26.3 15.2 19.9 42.1 20.3 40.0 28.8 30.4
LLaVA-v1.6-vicuna-13b 0.3B 23.3 19.6 24.5 23.1 36.7 26.9 47.5 28.5 33.2
Fuyu-8b - 21.7 22.1 27.3 23.3 27.3 24.4 27.1 24.9 25.1
Mini-Gemini-8b 0.3B 37.3 29.6 31.8 30.4 51.5 30.6 56.3 36.1 41.4
LLaVA-v1.5-7b 0.3B 18.0 25.0 15.8 19.7 41.5 19.7 35.0 25.7 28.4
LLaVA-v1.6-vicuna-7b 0.3B 21.3 22.9 18.2 20.5 36.5 19.4 32.9 26.9 31.5
OSS LLaVA-v1.6-mistral-7b 0.3B 24.3 25.8 24.5 24.9 38.5 24.2 36.7 32.1 28.9
(Small) Deepseek-VL-7b 0.38B 32.3 29.2 22.1 28.3 50.4 24.4 54.2 32.4 37.5
Yi-VL-6b 0.6B 25.2 35.5 26.2 28.8 35.6 29.0 54.5 30.8 34.4
Qwen-VL 1.9B 18.7 23.8 25.2 22.5 41.0 27.5 42.5 30.1 32.1
CogVLM2-LLaMA3-Chinese 10B 29.7 21.7 29.7 26.5 54.8 27.2 37.9 40.3 38.7
We report the accuracy of each type of question for the 24 models alongside the average human performance for each cognitive ability in Table. 1. We categorize the models into groups by their types, where open-source (OSS) MLLMs are grouped according to the size of their LLMs. It’s observed that even the most advanced MLLMs only marginally meet the passing line (60) for overall accuracy, e.g., Gemini-1.5-Pro (62.4) / GPT-4o (59.8) vs human (76.9). Notably, these models excel in domains related to verbal skills and knowledge, such as Gc and Grw. This success can likely be attributed to the powerful language capabilities inherent in large language models, bolstered by their extensive training datasets.
However, a significant performance gap remains between MLLMs and humans in areas like Visual-Spatial Abilities (Gv) and Fluid Reasoning (Gf). This is particularly evident in the Visual-Spatial Abilities domain, where all models lag considerably behind human capabilities, e.g., Gemini-1.5-Pro (53.8) vs human (81.1). This underscores a substantial opportunity for advancements in the visual aspects of MLLMs. See supplementary for case studies. Furthermore, our findings also highlight a pronounced deficiency in the Fluid Reasoning (Gf) capability among all MLLMs, particularly in tasks involving Induction (I) and Quantitative Reasoning (RQ). However, it is surprising to note that in the domain of Deductive Reasoning (RG), the most advanced MLLMs, such as GPT-4o, are approaching the average human level with scores of 59.2 compared to 60.0 for human participants. This might be attributed to the strategy they use synthetic reasoning data to enhance such ability [10].
Overall, MLLMs perform well in crystallized intelligence (Gc), possibly owing to their extensive training data, while the most advanced MLLMs still have a large gap with humans in fluid intelligence. This proves that our benchmark M3GIA can measure the difference between crystallized intelligence and fluid intelligence of MLLMs from a cognitive perspective, which is the key difference between M3GIA and other benchmarks.
Figure 4: (Left) The correlation matrix of MLLMs’ accuracy scores across various cognitive factors. It indicates a notable correlation in their performance across these dimensions. (Right) We find that the GIA scores of MLLMs, calculated using the CFA model constructed from human data, demonstrate high explanatory power for their overall test performance ($R^{2}\geq 0.93$).
Winner Takes All.
More importantly, our finding reveals an intriguing Winner Takes All phenomenon that merits further attention beyond the initial observations. Specifically, we noted a consistent trend within each group of models where proficiency in one cognitive domain often translates into superior performance across other domains as well in Table 1. In particular, despite the diversity in score distribution among different abilities, there is a noteworthy pattern: the models achieving the top and second-best scores across various cognitive abilities are predominantly the same two models within each group.
This shows an interesting consistency to the pattern observed in human intelligence which empirically suggests the existence of General Intelligence Ability (see Sec. 2). Therefore, it offers compelling evidence that general intelligence ability, also identified as the general factor of intelligence (“g”) at the stratum III of the CHC model, has also emerged in large models. Furthermore, it suggests that as MLLMs evolve towards more comprehensive cognitive processes, they too demonstrate a foundational GIA factor that simultaneously governs a variety of cognitive abilities.
4.2 Multilingual GIA Scores
By collecting a large amount of testing data from human subjects, we adoptted CFA (Confirmatory Factor Analysis) model to calculate the GIA scores which can reflect comprehensive intelligence factors. Since the questions for each language are not exactly the same, we need to establish a separate CFA model for each language. To our surprise, the model built with human data showed high explanatory validity for the test results of MLLMs (cor > 0.93). This indicates, to some extent, that the cognitive structure of MLLMs indeed shows similarities to humans. We report the GIA scores of each language for some MLLMs of different sizes in Table. 2 and Fig. 5. It’s observed that the current state-of-the-art MLLMs have reached the minimum level within the human subjects’ confidence interval in English. However, these MLLMs still exhibit a significant performance gap compared to humans in other languages.
Table 2: The GIA scores of different models accross the six languages. The best in bold and the second-best underlined. The lower the better.
Models General Intelligence Abilitiy (GIA) Normalized GIA Scores
En Ch Fr Sp Pt Ko En Ch Fr Sp Pt Ko
Human 16.01 16.69 19.52 16.22 16.00 18.05 100.0 100.0 100.0 100.0 100.0 100.0
GPT-4o 13.85 11.46 12.37 13.12 12.80 13.01 86.5 68.7 63.3 80.9 80.0 72.1
GPT-4v 12.61 10.95 13.83 14.04 12.12 12.25 78.8 65.6 70.8 86.5 75.8 67.9
LLaVA-1.6-34b 11.47 9.25 11.35 7.96 10.67 9.04 71.6 55.4 58.1 49.1 66.7 50.1
LLaVA-1.6-13b 6.96 6.89 8.71 7.75 6.94 7.75 43.5 41.3 44.6 47.8 43.4 42.9
LLaVA-1.6-7b 6.75 5.99 7.67 6.74 6.01 5.93 42.1 35.9 39.3 41.5 37.6 32.9
Mini-Gemini-34b 11.00 9.96 12.75 11.52 9.45 10.69 68.7 59.7 65.3 71.0 59.1 59.2
Mini-Gemini-13b 8.68 7.76 8.65 7.73 7.10 7.98 54.2 46.5 44.3 47.7 44.4 44.2
Mini-Gemini-8b 9.32 8.11 11.25 9.76 8.99 7.08 58.2 48.6 57.6 60.2 56.2 39.3
Qwen-72b 11.68 10.75 10.20 10.50 9.71 9.76 72.9 64.4 52.2 64.7 60.7 54.1
Qwen-32b 10.58 9.79 9.62 10.11 9.25 9.18 66.1 58.7 49.3 62.3 57.8 50.9
Qwen-14b 8.46 8.76 9.15 8.49 8.79 8.32 52.8 52.5 46.9 52.4 54.9 46.1
Qwen-7b 8.56 8.93 8.98 8.42 8.77 8.41 53.4 53.5 46.0 51.9 54.8 46.6
Qwen-1.8b 7.34 6.56 8.01 7.37 6.49 6.48 45.8 39.3 41.0 45.5 40.6 35.9
Figure 5: The GIA scores of some representative MLLMs across the six languages. (Left) In each language, we set the average human score as 100 and use this to normalize the scores of different models. Consequently, the GIA score of each language becomes comparable. (Right) We take English as an example and visualized the GIA performance of various models and compared it with the human level.
To further investigate the influence of LLM size to the GIA score, we conducted an ablation study with the Qwen series. In order to strictly control variables like different training data and ViT components. we trained the models by ourselves using the same training data for pretraining and fine-tuning and we also use the same ViT component (CLIP-ViT-L-14) in the series. Overall, the GIA scores of the models increase with the rise in LLM parameters. However, we observed a somewhat counterintuitive phenomenon. There is often no improvement in cognitive abilities from 7B to 13B, and seems to be a emerging point of General Intelligence Ability between 13b and 34b.
5 Conclusion
This paper has presented M3GIA, the first evaluation benchmark that comprehensively evaluate the cognitive abilities of MLLMs under the theoretical umbrella of the well-recognized Cattell-Horn-Carroll (CHC) Model of Intelligence. Based on the CHC theory, we identified five key cognitive factors for current MLLMs: Fluid reasoning (Gf), Comprehension-Knowledge (Gc), Visual processing (Gv), Reading and Writing (Grw), Quantitative knowledge (Gq), and designed five broad types of questions to measure them. In order to meet the pressing need for multilingual assessment, our evaluation data spans across six languages and are collected from native language sources, including English, Chinese, French, Spanish, Portuguese and Korean. We conducted a series of experiments to comparative analyze the cognitive abilities of various MLLMs against human performance, and discussed how factors like the size of LLM component impact cognitive abilities.
6 Limitations and Discussions
This version of M3GIA does not include all the broad cognitive factors of the CHC model, such as auditory processing (Ga), olfactory processing (Go), etc. As advancements in MLLMs incorporate a broader range of modalities, more factors from the CHC framework can be integrated, ensuring that M3GIA remains at the forefront of evaluating future generations of multimodal models.